1. If y = (x^2 + R^2)^(3/2) where R is a positive constant then find dy/dx.

Grade:11

1 Answers

AskiitianExpert Shine
14 years ago

Hi

y = (x^2 + R^2)^3/2
 
dy/dx = d ((x^2 + R^2)^3/2) / dx
          = 3/2 ((x^2 + R^2)^1/2) x 2x
          = 3x(x^2 + R^2)^1/2